CV-230, AOC-223
Dial-in: 434-972-7268

Attendees: Carol, Kimberly, Scott, Rachel, Anthony, Aaron, Adam, Dongchan, Jonathan, Lyndele Gerald, James at HIA

1. News and overview
  • Cycle 1 schedule -
    • proposed scheduled forwarded by John Hibbard - July 12th deadline
    • telecons continuing on Tuesdays mornings to work on documentation
    • call not until April
    • internal info only- JAO has not made this announcement
    • technical reviews will be done afterwards on the selected proposals
    • G. Mathys drafting the review process and will submit to ARC managers for review
    • Is there a list somewhere that shows the progress of work on each document? No, currently being managed by Working Group

Check to see if Working Group telecon can be held M, W or F early morning. If so, Gerald could participate
  • Jonathon Tan has his data and Dan Marrone's data are here.

2. Coordination with other groups and activities
  • Site, CSV, SCO
  • DSG & computing
  • JSG
  • Tiger Team
    • met on Tuesday; now links from main CASA page to packages.
  • Software

3. Main items

4. USG task reports & new business
  • HIA (Gerald)
    • Press release in mid-Feb about completion of Band 3 cartridge'
    • Mark M. attending celebration in Feb
    • Future of m.sub-m astronomy in Canada. Rachel will be attending, Feb 17th in Toronto, Dunlap Institute
  • Taiwan and EA Liaison (Aaron)
    • Chen Fai - sent email to begin dialogue about collaborations. Carol will forward emails about working on next meeting.
  • Documentation (Jim, Jonathan)
    • Trying to talk with Phil Appleton; hoping to talk about lessons learned.
    • New Herschel webinars coming up
    • Synthesis Imaging - online mm course and material for Summer Institute may be similar? Adam: hands-on OT portion and EVLA session and Data Reduction. Adam and Tony are going. Anyone else going? Possibly Kimberly Scott. May 29-June 8. Talks will include ALMA and EVLA.
    • Jonathan and Jim meeting tomorrow to talk more about mm course
  • Helpdesk (Tony)
    • Int'l working group meeting 1/25. Change request issues that are not streamlined - some requests are going all the way to Thijs. Need to better understand flow.
    • Attempting to increase visibility and encourage use of Knowledge Base articles.
  • PI Support & Visitors (Scott)
  • Development of outreach materials (Carol)
  • Workshops, Schools, Meetings (Tony, Adam)
    • Logistics meeting with IT to prepare for JETS 2012; all under control.
  • Tutorials (Carol)
  • Science Web (Aaron)
    • Web pages 'getting stale' (Fred Lo) - simplify NAASC page
    • Aaron will send around document outlining ideas
    • DA in Green Bank will be a web page guru

5. AOB

Current USG Focus Items
  • Batch 3 and batch 5 SBs
    • Cone searches have not been progressing; reminded via Jira ticket that project can't move forward
    • John found a problem with blocks - working with Harvey; taken care of.
  • Have been promised extra 100GB
  • 211 (Visitor Office) will get three new workstations, so all six desks will have workstations.
  • Feb 11 major downtime
  • Cycle 1 documents
  • Simulations and science examples

-- LyndeleVonSchill - 2012-01-26
Topic revision: r1 - 2012-01-26, LyndeleVonSchill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback